business finance

Carson Trucking is considering whether to expand its regional service center in Mohab, UT.  The expansion requires the expenditure of

$

on new service equipment and would generate annual net cash inflows from reduced costs of operations equal to

$

per year for each of the next

years.  In year

the firm will also get back a cash flow equal to the salvage value of the equipment, which is valued at

$

million.  Thus, in year

the investment cash inflow totals

$.

Calculate the project's NPV using a discount rate of

percent.
Answer #1

Net present value = present value of cash inflows - present value of cash outflows

present value of cash inflows = 3,000,000 / ( 1 + 0.07)1+ 3,000,000 / ( 1 + 0.07)2+ 3,000,000 / ( 1 + 0.07)3+ 3,000,000 / ( 1 + 0.07)4+ 3,000,000 / ( 1 + 0.07)5+ 3,000,000 / ( 1 + 0.07)6+ 4,100,000 / ( 1 + 0.07)1

present value of cash inflows = 2,803,738.318 + 2,620,316.185 + 2,448,893.631 + 2,288,685.636 + 2,138,958.538 + 1,999,026.671 + 2,553,273.942

present value of cash inflows = 16,852,892.92

NPV = 16,852,892.92 - 10,500,000

NPV = $6,352,892.921


answered by: ANURANJAN SARSAM
Answer #2

SOLUTION :


Discount rate, r = 6% = 0.06 

=> 1 + r = 1.06


We know that :


NPV 

= - Initial cost + PV of future cash flows discounted @ 6%.

= - 9000000 + 3500000(1.06^7 - 1)/(0.06 * 1.06^7) + 900000/1.06^7

= 11136886.45 ($) 


Investment cash inflow is not considered as it is not generated from the project.


So, NPV of the project is = 11,136,886 approx.  ($) (ANSWER)


answered by: Tulsiram Garg